OpenStudy (anonymous):

Determine the right end-behavior of the function h(t)=3[(t+1)/t]^t e3 0 3e infinity

zepdrix (zepdrix):

So what's happening for large values of t?\[\Large\bf\sf \lim_{t\to\infty}3\left(\frac{t+1}{t}\right)^t\] Hmm

zepdrix (zepdrix):

If you split up the fractions inside, \[\Large\bf\sf \lim_{t\to\infty}3\left(1+\frac{1}{t}\right)^t\]You should be able to relate this to the limit definition of e. Does it look familiar?